Debra borrowed $8000 at a rate of 11%, compounded semiannually. Assuming she makes no payments, how much will she owe after 5 years. do not round any intermediate computations and round your answer to the nearest cent
Answer by lwsshak3(11628)

Debra borrowed $8000 at a rate of 11%, compounded semiannually. Assuming she makes no payments, how much will she owe after 5 years. do not round any intermediate computations and round your answer to the nearest cent.
***
Compound interest formula: A=P(1+i)^n, P=initial investment, i=interest rate per period, n=number of periods.
For given problem:
P=8000
i=.11/2=.055
n=5*2=10
A=amt after n periods
..
A=8000(1+.055)^10
A=8000(1.055)^10
A=$13,665.16
After 5 years, Brenda will owe $13,665.16